### v1

@@ Abstract
Few-Shot Learning, Few-Shot Object Detection, Deep Learning Model Generalization, Feature Normalization, Low-Data Machine Learning, Cross-Domain Benchmark, Bangladeshi Traffic Sign Detection
@@ I Introduction
Over the years, traffic sign detection systems have evolved from manually engineered features to Deep Learning (DL)-based methods, particularly Convolutional Neural Networks (CNNs) [6]. These methods are generally categorized into two categories: one- and two-stage detectors [7]. One-stage detectors such as YOLO [8] treat sign detection as a regression problem. Here, a single convolutional network simultaneously predicts multiple bounding boxes around the traffic sign along with the class probabilities for those boxes, offering faster processing but less accuracy [9, 10]. In contrast, two-stage detectors, such as Faster R-CNN [11], first generate region proposals around the signs and then classify them, leading to higher accuracy but slower inference.
Despite the emergence of DL-based methods that have introduced significant progress in traffic sign detection [12, 13], these methods require extensive annotated training data to achieve optimal performance. The publicly available traffic sign datasets often focus on specific countries, with limited traffic sign categories and low variability, which limits the effectiveness of these methods. As a result, a model trained on one country’s dataset might not perform well on another without being fine-tuned on a sufficient number of target samples. Moreover, despite the advancements in hardware technology and the generation of massive datasets that have propelled the DL-based approaches towards human-like capabilities, they are still unable to learn effectively from only a few samples [14]. Unlike humans, who can quickly adapt to new tasks using prior knowledge, DL models require extensive data to perform effectively [15]. This disparity in learning efficiency is particularly challenging in the context of traffic sign detection in underrepresented regions, such as Bangladesh, where annotated samples are scarce. In this connection, the recent family of Few-Shot Learning (FSL)-based models can play an instrumental role with its remarkable ability to learn with limited data and has already proved its worth in a diversified arena of computer vision tasks [16].
@@ III Proposed Methodology
Our proposed framework is built upon Faster R-CNN, pre-trained on the MS COCO dataset [32], due to its established robustness in object detection tasks [33]. Recognizing the challenges inherent in few-shot traffic sign detection, we introduced several key modifications to adapt Faster R-CNN for this specialized task. These modifications are designed to enhance the model’s ability to generalize from a small number of examples, a crucial capability for effective few-shot learning.
We employ a few key strategies: keeping the entire network unfrozen to allow all layers to adapt to the new data and creating a pseudo-support set through data augmentation, which increases the diversity of the training data and enhances the model’s resilience to real-world variations. Again, we replace the original classifier with a cosine similarity-based classifier to enable embedding normalization that enhances generalization from limited training samples by focusing on intrinsic relationships between instances and classes. This approach mitigates high intra-class variance, which is particularly challenging in few-shot scenarios. Furthermore, we fine-tune the entire network end-to-end using the MTSD dataset to ensure that the learned features are specifically tailored to the characteristics of traffic signs, which differ from those of general objects in the COCO dataset. Finally, we evaluate the performance of FUSED-Net on the target dataset across various few-shot scenarios (1, 3, 5, and 10-shot) to ensure that it is well-suited to the specific challenges of traffic sign detection in a data-scarce environment.
@@ III-A Baseline Faster-RCNN
The object detection process traditionally follows three main steps: generating region proposals that might contain objects, extracting feature vectors for these proposals, and classifying each feature vector as either background or a specific object class. Early models like RCNN implemented this pipeline by utilizing algorithms like selective search [36] to generate region proposals. For feature extraction, RCNN replaced traditional descriptors like histograms of oriented gradients [37] with CNNs, dramatically improving accuracy. However, RCNN used a pre-trained Support Vector Machine [38] for classification, which required extracting and caching features from every region proposal separately—a process that was computationally expensive and slow.
Fast R-CNN [35] addressed these inefficiencies by introducing the Region of Interest (ROI) pooling layer, which allowed for the extraction of fixed-size feature vectors directly from a shared feature map. This innovation enabled all region proposals to share computation, significantly speeding up the process. Furthermore, Fast R-CNN eliminated the need for caching features on disk, as it could process all proposals within a single forward pass through the network, thereby streamlining the object detection pipeline.
@@ III-B Unfrozen Parameters
When training deep learning models, it is common practice to ‘freeze’ certain layers, especially when fine-tuning a pre-trained model on a new task [45]. Freezing layers means that the weights of those layers are kept constant during the training process, and only the weights of the unfrozen layers are updated. This technique is particularly useful when the new task is similar to the original task the model was trained on, as it helps retain the learned features while allowing the model to adapt to the new task with minimal computational overhead [34, 46, 47].
However, in the context of cross-domain tasks, few-shot traffic sign detection in particular, the target domain (e.g., traffic signs from a specific region or country) can be significantly different from the base domain (e.g., general object detection on datasets like MS COCO) representing unique environmental challenges [39, 48, 49]. In such cases, freezing layers can be counterproductive, as it limits the model’s ability to learn new features that are crucial for accurately detecting and recognizing objects in the target domain. The frozen layers may still carry biases and features specific to the base domain, leading to suboptimal performance when the model is applied to the target domain [50].
To overcome this limitation, we chose to keep the entire architecture unfrozen during the fine-tuning stage. This means that all components of our detector, including the Backbone, Region Proposal Network (RPN), Feature Pyramid Network (FPN), and fully connected layers, are updated during training. By allowing every part of the network to adapt, we ensure that the model can learn new, domain-specific features that are essential for detecting traffic signs in different environments. This approach contrasts with the conventional method of freezing certain layers and allows our model to better handle the significant differences between the base and target domains. As a result, our model demonstrates improved adaptability and accuracy in the target domain.
@@ III-C Pseudo Support Sets (PSS)
In FSOD, the limited availability of labeled training data presents a significant challenge in developing robust models. To address this, data augmentation is a common technique used to enhance the diversity and variability of the training data. In our methodology, we implemented a data augmentation strategy during the fine-tuning stage by generating a ‘Pseudo-Support Set (PSS)’ by augmenting the original support set. The creation of this pseudo-support set involves applying random Color Jitter to each sample in the original support set. Color Jitter is a data augmentation technique that randomly alters the brightness, contrast, saturation, and hue of an image. These specific transformations were chosen because they closely mimic the kinds of variations that are commonly encountered in real-world scenarios, such as changes in lighting conditions, different times of the day, weather conditions, and even camera settings. Brightness adjustment simulates conditions ranging from very bright, sunny environments to dim, overcast scenarios. Contrast alteration affects the differentiation between light and dark areas in an image, allowing the model to recognize objects in both high-contrast situations like strong shadows, and low-contrast environments like foggy weather. Saturation changes the intensity of colors, helping the model learn to detect traffic signs that might appear more muted or vividly colored due to different camera sensors or environmental conditions. Hue adjustment shifts the overall color balance of the image, which can account for the color variances caused by different lighting sources (e.g., fluorescent vs. natural light).
By creating PSS through this augmentation process, we introduce a broader range of visual characteristics into the training data. The PSS, which combines the original samples with their augmented versions, provides the model with a richer and more varied dataset. This helps the model learn to recognize patterns under diverse conditions, improving its generalization ability and reducing the risk of overfitting to the limited original data.
This approach also enhances the model’s robustness, making it less sensitive to slight changes in input data that could lead to misclassification. By training on the pseudo-support set, the model becomes better equipped to handle the unpredictable variations it will encounter in real-world traffic sign detection tasks, ensuring more reliable and accurate performance across different environments.
@@ III-D Embedding Normalization (EN)
Embedding Normalization (EN) is a key component of our proposed method, especially in the context of few-shot learning, where the challenge of high intra-class variance becomes particularly pronounced [20]. In few-shot scenarios, the model is trained on a limited number of examples, which can lead to significant variation in feature representations within the same class. This variance often results in a model that is overly sensitive to outliers, thereby compromising its generalization ability. To mitigate this issue, we implemented EN to standardize the feature representations of individual instances. This standardization process ensures that the features extracted from each instance are on a comparable scale, regardless of the inherent differences in the dataset. By reducing the intra-class variance, the model becomes more robust to variations within the same class, which is essential in few-shot learning where such differences can otherwise lead to misclassifications. The normalization technique plays a pivotal role in aligning the features more closely, making it easier for the classifier to distinguish between different classes even when only a few examples are available for training.
To incorporate EN, we integrated a cosine similarity-based classifier into our model. This classifier focuses on the angular relationships between feature vectors rather than their magnitudes. By emphasizing the direction rather than the length of the vectors, the cosine similarity-based classifier inherently reduces the influence of large magnitude differences, which could otherwise skew the classification results. This characteristic is particularly beneficial in few-shot learning, where the limited data can result in embedding vectors with varying magnitudes that do not necessarily reflect meaningful differences between instances.
The incorporation of EN improves the model’s ability to generalize from a few examples. It ensures that the model’s decision-making process is less influenced by outlier instances and more focused on the intrinsic relationships between instances and classes. In the context of few-shot traffic sign detection, where the training data is sparse and diverse, this approach helps the model to better capture the essential features of traffic signs, leading to more accurate and reliable classification. This enhancement is particularly important for our application, where the model needs to perform well despite the limited and varied nature of the available data.
@@
This approach emphasizes the alignment of embedding vectors, making the classifier more robust to variations in embedding magnitude and improving the model’s ability to generalize to new, unseen classes.
@@ III-E Domain Adaptation
In our approach, we employ domain adaptation by training our model on traffic signs from different countries, thereby exposing it to a diverse set of visual features and styles. This diverse training data helps minimize the distributional discrepancy between the base domain (e.g., traffic signs in the training dataset) and the target domain (e.g., traffic signs from another country with limited data). By leveraging this approach, our model becomes more generalized and robust, capable of accurately detecting and recognizing traffic signs across different domains. The effectiveness of this strategy is reflected in the improved performance of our model, which we will discuss in the results section.
@@ IV Results & Discussion
In this section, we present a comprehensive analysis of our proposed FUSED-Net framework, focusing on the datasets used, implementation details, ablation study, error analysis, and performance comparisons with state-of-the-art models. We introduce the Merged Traffic Sign Detection Dataset (MTSD) and the Bangladeshi Traffic Sign Detection Dataset (BDTSD) to highlight the diverse and challenging conditions under which our model was trained and tested. We also discuss our implementation within the Detectron2 framework, detailing the hyperparameters and experimental environment. We follow that up with an ablation study showing the incremental benefits of each component in our model, while performance comparisons underscore FUSED-Net’s superiority over other models in various few-shot traffic sign detection scenarios. Finally, we provide qualitative and cross-domain performance analyses, demonstrating the robustness and adaptability of FUSED-Net in real-world and cross-domain settings.
IV-A Datasets
@@
IV-A1 Merged Traffic Sign Detection Dataset (MTSD)
One of our primary objectives in constructing MTSD was to maximize the variation in the samples. To achieve this, we selected datasets from two different continents: Europe (GTSDB, DFG) and America (LISA), where the traffic signs differ significantly due to variations in regulatory standards, design philosophies, and cultural norms. For example, U.S. speed limit signs are rectangular with black text on a white background (Fig. 3c), while European speed limit signs are circular with a red border and black numbers (Fig. 3a). Additionally, European signs often utilize symbols and pictograms without text, enhancing accessibility across different languages, whereas American signs frequently rely on text, which can be less intuitive at times. To add to the complexity, as shown in Fig. 3b, in some European countries, traffic signs may appear on the left side of the road and can consist of multiple signs.
The GTSDB and LISA datasets feature German and American traffic signs, respectively, covering a total of 90 categories. Both datasets are widely used in the field of traffic sign detection. The DFG dataset comprises 200 categories and is, to our knowledge, the only traffic sign detection dataset with such an extensive number of classes.
In total, we compiled 288 categories in MTSD from the aforementioned datasets. Although we had 14,416 samples, the presence of multiple traffic signs within a single sample led to a total of 23,276 annotations. The “roundabout” and “stop” categories appeared in multiple datasets. Since different datasets contained varying samples for the same category, which could potentially confuse our architecture, we excluded these two categories from our compilation. Given that the original Faster R-CNN was trained on Pascal VOC, which contains 24,640 annotations of 9,963 sample images, the MTSD dataset should be sufficient for base training any architecture using Faster R-CNN as its base detector. Therefore, we did not include any datasets beyond these three.
In our proposed framework that incorporates domain adaptation, MTSD is utilized for the initial training of our architecture. This approach enables the system to better comprehend traffic sign characteristics, enhancing its capability to detect unfamiliar traffic signs during the fine-tuning stage.
IV-A2 Bangladeshi Traffic Sign Detection Dataset (BDTSD)
@@
IV-A3 Cross-Domain Few-Shot Object Detection (CD-FSOD) Benchmark

### v2

@@ Abstract
Few-Shot Learning, Unfrozen Parameters, Pseudo-Support Sets, Embedding Normalization, Domain Adaptation, Cross-Domain Benchmark, Bangladeshi Traffic Sign Detection,
@@ I Introduction
Over the years, traffic sign detection systems have evolved from manually engineered features to Deep Learning (DL)-based methods, particularly Convolutional Neural Networks (CNNs) [6]. These methods are generally categorized into two categories: one- and two-stage detectors [7]. One-stage detectors such as YOLO [8, 9] and DETR [10] treat sign detection as a regression problem. Here, a single convolutional network simultaneously predicts multiple bounding boxes around the traffic sign along with the class probabilities for those boxes, offering faster processing but less accuracy [11, 12]. In contrast, two-stage detectors, such as Faster R-CNN [13], first generate region proposals around the signs and then classify them, leading to higher accuracy but slower inference.
Despite the emergence of DL-based methods that have introduced significant progress in traffic sign detection [14, 15], these methods require extensive annotated training data to achieve optimal performance. However, publicly available datasets, such as TT100K [14], primarily represents traffic signs from specific regions like China, which are significantly different from those in underdeveloped or developing countries, such as Bangladesh. In these countries, the condition of traffic signs is often compromised due to limited maintenance, occlusion, or wear and tear, introducing additional challenges that are not well-represented in such datasets. Consequently, a model trained on one country’s dataset might not perform well on another without being fine-tuned on a sufficient number of target samples. In this context, collecting large-scale annotated datasets for traffic signs across every country is impractical due to the diversity of sign designs and environmental conditions.
Even with the advancements in hardware technology and the generation of massive datasets that have propelled the DL-based approaches towards human-like capabilities, they are still unable to learn effectively from only a few samples [16]. Unlike humans, who can quickly adapt to new tasks using prior knowledge, DL models require extensive data to perform effectively [17]. This disparity in learning efficiency is particularly challenging in the context of traffic sign detection in underrepresented regions, such as Bangladesh, where annotated samples are scarce. In this connection, the recent family of Few-Shot Learning (FSL)-based models can play an instrumental role with its remarkable ability to learn with limited data and has already proved its worth in a diversified arena of computer vision tasks [18], making them an ideal choice for this problem domain.
@@ III Proposed Methodology
Our proposed framework is built upon Faster R-CNN, pre-trained on the MS COCO dataset [34], due to its established robustness in object detection tasks [35]. Recognizing the challenges inherent in few-shot traffic sign detection, we introduced several key modifications to enhance the model’s ability to generalize from a small number of examples. First, we keep the entire network unfrozen during training, allowing all layers to adapt to the new data and enabling more effective learning of domain-specific features. Second, we augment the training data by creating a pseudo-support set, which combines original samples with their augmented versions, increasing diversity and improving the model’s resilience to real-world variations. To further mitigate the high intra-class variance inherent in few-shot scenarios, we replace the original classifier with a cosine similarity-based classifier. This replacement incorporates embedding normalization, focusing on the intrinsic relationships between instances and classes to ensure better generalization from sparse training samples.
Additionally, we fine-tune the entire network end-to-end using the MTSD dataset, ensuring that the features learned are specifically tailored to traffic signs, which present unique characteristics distinct from the general objects in COCO. The effectiveness of these strategies is evaluated by testing FUSED-Net on the target dataset across various few-shot scenarios (1, 3, 5, and 10-shot), demonstrating its suitability for the specialized task of traffic sign detection in data-scarce environments. The overview of the FUSED-Net Architecture is shown in Fig. 2.
@@ III-A Baseline Faster-RCNN
The object detection process traditionally involves three steps: generating region proposals, extracting feature vectors, and classifying them. Early models like RCNN used Selective Search [38] for proposal generation and CNNs for feature extraction, improving accuracy over traditional descriptors. However, RCNN relied on a pre-trained Support Vector Machine [39] for classification, which was computationally expensive due to the need for caching features from each region proposal. Fast R-CNN improved upon this by introducing the ROI pooling layer, enabling the extraction of fixed-size feature vectors from a shared feature map, allowing proposals to share computation. This innovation sped up the process and eliminated the need for caching, significantly streamlining the detection pipeline.
@@ III-B Unfrozen Parameters
In deep learning, “freezing” certain layers during training is common, especially when fine-tuning pre-trained models on a new task [46]. Freezing keeps the weights of those layers fixed, updating only the unfrozen layers. This helps retain learned features while minimizing computational overhead, particularly when the new task is similar to the original task [36, 47, 48]. However, in cross-domain tasks like few-shot traffic sign detection, where the target domain (e.g., region-specific traffic signs) differs significantly from the base domain (e.g., general object detection), freezing layers can hinder performance [40, 49, 50]. The model may struggle to learn new features essential to the target domain, as frozen layers may carry biases and features from the base domain, reducing detection accuracy [51].
To overcome this limitation, we chose to keep the entire architecture unfrozen during the fine-tuning stage. This means that all components of our detector, including the Backbone, Region Proposal Network (RPN), Feature Pyramid Network (FPN), and fully connected layers, are updated during training. Given the network with layers $L_{1},L_{2},\dots,L_{n}$, unfrozen training updates the weights of all the layers such that $\nabla W_{L_{i}}\neq 0$ for $i=1,\dots,n$. By allowing every part of the network to adapt, we ensure that the model can learn new, domain-specific features that are essential for detecting traffic signs in different environments. This approach contrasts with the conventional method of freezing certain layers and allows our model to better handle the significant differences between the base and target domains. As a result, our model demonstrates improved adaptability and accuracy in the target domain.
@@ III-C Pseudo Support Sets (PSS)
In FSOD, the limited availability of labeled training data presents a significant challenge in developing robust models. To address this, data augmentation is a common technique used to enhance the diversity and variability of the training data. In our methodology, we implemented a data augmentation strategy during the fine-tuning stage by generating a ‘Pseudo-Support Set (PSS)’ by augmenting the original support set. Mathematically, let $X_{s}$ represent the original support set and $T$ represent the transformation applied. The augmented support set $X_{s}^{\prime}$ is then defined as:
$X_{s}^{\prime}=X_{s}|T(X_{s})$
The creation of this pseudo-support set involves applying random Color Jitter to each sample in the original support set. Here, we chose to use only Color Jitter for augmentation, guided by [52], who showed that stronger color jittering is effective in tasks requiring visual representations.
Color Jitter is a data augmentation technique that randomly alters the brightness, contrast, saturation, and hue of an image. These specific transformations were chosen because they closely mimic the kinds of variations that are commonly encountered in real-world scenarios, such as changes in lighting conditions, different times of the day, weather conditions, and even camera settings. Brightness adjustment simulates conditions ranging from very bright, sunny environments to dim, overcast scenarios. Contrast alteration affects the differentiation between light and dark areas in an image, allowing the model to recognize objects in both high-contrast situations like strong shadows, and low-contrast environments like foggy weather. Saturation changes the intensity of colors, helping the model learn to detect traffic signs that might appear more muted or vividly colored due to different camera sensors or environmental conditions. Hue adjustment shifts the overall color balance of the image, which can account for the color variances caused by different lighting sources (e.g., fluorescent vs. natural light).
The creation of PSS by supplementing original samples with their augmented versions introduces a broader range of visual characteristics into the training data. This enriched dataset helps the model recognize patterns under diverse conditions, improving generalization and reducing overfitting. Additionally, training on the PSS enhances the model’s robustness, making it less sensitive to input variations and better equipped to handle the unpredictable changes encountered in real-world traffic sign detection, leading to more reliable and accurate performance.
@@ III-D Embedding Normalization (EN)
Embedding Normalization (EN) is crucial in our method, particularly considering our few-shot learning scenario, where high intra-class variance is a challenge [22]. In few-shot scenarios, limited training examples lead to significant variations in feature representations, making the model overly sensitive to outliers and reducing generalization. To address this, EN standardizes feature representations, ensuring that all instances are on a comparable scale, which reduces intra-class variance and enhances robustness.
To incorporate EN, we integrate a cosine similarity-based classifier into FUSED-Net, which focuses on angular relationships between feature vectors, rather than magnitudes. This approach helps mitigate the impact of varying vector lengths, ensuring more accurate detection, especially in few-shot learning where limited data can lead to inconsistent feature magnitudes. EN improves generalization by reducing the influence of outliers and aligning features more effectively, allowing the classifier to distinguish between classes with minimal data. This is particularly beneficial for our application domain, few-shot traffic sign detection, where sparse and diverse data is common.
@@
The range of cosine similarity, $sim_{(x_{i},w_{j})}$ varies between $-1$ to $+1$. A value of $+1$ indicates the vectors $\hat{x}_{i}$ and $\hat{w}_{j}$ are perfectly aligned in the same direction, implying a high degree of similarity between the instance and the class. A value of $-1$ suggests that the vectors are diametrically opposed, indicating dissimilarity. A value near 00 indicates orthogonality, meaning there is no discernible similarity between the two vectors. The value of $\gamma$ is determined empirically.
@@
By considering cosine similarity, we ensure that the model is invariant to vector magnitudes and instead learns to prioritize the orientation of the feature vectors, which is particularly useful for few-shot learning with limited data.
@@ III-E Domain Adaptation
In our approach, we employ domain adaptation by training our model on traffic signs from different countries, thereby exposing it to a diverse set of visual features and styles. This diverse training data helps minimize the distributional discrepancy between the base domain (e.g., traffic signs in the training dataset) and the target domain (e.g., traffic signs from another country with limited data). Let $P_{S}(X,Y)$ and $P_{T}(X,Y)$ represent the source and the target distributions, respectively. By training on diverse traffic sign datasets from multiple countries, our model minimizes the discrepancy $\mathcal{D}(P_{S},P_{T})$. By leveraging this approach, our model becomes more generalized and robust, capable of accurately detecting and recognizing traffic signs across different domains. The effectiveness of this strategy is reflected in the improved performance of our model, which we will discuss in the Results section.
@@ IV Results & Discussion
In this section, we present a comprehensive analysis of our proposed FUSED-Net framework. We begin with the experimental setup, which introduces the datasets used, namely Merged Traffic Sign Detection Dataset, Bangladeshi Traffic Sign Detection Dataset, and Cross-Domain Few-Shot Object Detection Benchmark, showcasing the diverse and challenging conditions under which our model was trained and tested. Additionally, we provide the implementation details, including the hyperparameters and experimental environment within the Detectron2 framework. Next, we compare FUSED-Net with the state-of-the-art models in various few-shot traffic sign detection scenarios. This discussion also includes a per-category performance analysis, examining detection accuracy across the traffic sign categories in the query set, and the statistical significance of the reported performance based on multiple independent trials. Following this, we provide a deeper understanding of FUSED-Net’s functionality through an ablation study that highlights the incremental benefits of its components and a qualitative analysis illustrating its performance in challenging scenarios using representative examples. Finally, we assess FUSED-Net’s adaptability in generalized settings showcasing its results on the Cross-Domain Few-Shot Object Detection (CD-FSOD) benchmark, which underscores the model’s robustness and versatility in diverse real-world conditions.
IV-A Experimental Setup
@@
IV-A1 Datasets
Merged Traffic Sign Detection Dataset (MTSD)
The MTSD dataset is compiled from three established traffic sign detection datasets: The German Traffic Sign Detection Benchmark (GTSDB) [55], the LISA Traffic Sign Dataset (LISA) [56], and the DFG Traffic Sign Dataset (DFG) [15]. The GTSDB and LISA datasets feature German and American traffic signs, respectively, and together cover a total of 90 categories. Both of these datasets are widely used in the field of traffic sign detection. The DFG dataset, which is unique in its extensive collection, contains 200 categories, making it the largest among the three. In total, the MTSD dataset includes 272 categories, derived from these three sources.
Although MTSD contains 14,416 samples, the presence of multiple traffic signs within a single image leads to a total of 23,276 annotations. To prevent potential confusion in our architecture, we excluded two categories, “roundabout” and “stop”, which appeared in multiple datasets and had varying sample representations across them. This was done to avoid the possibility of conflicting data affecting model performance.
The dataset was designed with the aim of maximizing sample diversity. To this end, we selected traffic sign datasets from two different continents: Europe (GTSDB, DFG) and America (LISA). Traffic sign designs differ significantly between these regions due to variations in regulatory standards, design philosophies, and cultural norms. For example, U.S. speed limit signs are typically rectangular with black text on a white background (Fig. 3c), whereas European speed limit signs are circular with a red border and black numbers (Fig. 3a). Additionally, European traffic signs often feature symbols and pictograms without text, enhancing cross-linguistic accessibility, while American signs more commonly rely on text, which can be less intuitive. Moreover, in some European countries, traffic signs may be positioned on the left side of the road, and a single location may feature multiple signs (Fig. 3b).
To provide a clearer understanding of the dataset, we present the class distribution balance in MTSD in Fig. 4. The dataset exhibits a highly skewed class distribution: over 85% of the categories contain fewer than 100 samples, while some categories have more than 1000 samples. This imbalance could impact model performance in few-shot learning tasks, but it does not undermine the dataset’s primary purpose. The goal of MTSD is not to offer a balanced training set but to expose the model to a broad range of traffic sign variations from different countries. This diversity is intended to promote the learning of generalized features that can be applied across various visual styles and regulatory standards.
(a) Number of Samples Per Category
(b) Number of Annotations Per Category
Figure 4: The dataset statistics of MTSD. The sample and the annotation count per category is highly skewed considering more than 85% of the categories have less than 100 samples whereas the most frequent categories have more than 1000 samples.
In our proposed framework that incorporates domain adaptation, the MTSD dataset is used for the initial training of the model. This pretraining enables the system to better understand traffic sign characteristics, which enhances its ability to detect unfamiliar signs during the fine-tuning stage. Despite the dataset’s imbalanced nature, this approach helps the model generalize more effectively across domains, improving its performance on real-world traffic sign detection tasks.
Bangladeshi Traffic Sign Detection Dataset (BDTSD)
@@ IV-A1 Datasets
Cross-Domain Few-Shot Object Detection (CD-FSOD) Benchmark